Perpendicular lines AB and CD intersect at point E. If m∠AED = x + 20, what is the value of x?
my choices are as follows:
a) 160
b) 110
c) 90
or d) 70
i am alright in geometery but not great and i dont understand how to work this problem or i would find the answer myself. the problem is i cant find out how to work it at all neither on the internet or in my textbook!!!

Found 2 solutions by mananth, richard1234:
Answer by mananth(16946)

The lines are perpendicular.
so they intersect at right angles = 90 degrees
x+20 =90
x=90-20
x= 70 deg. option D

Answer by richard1234(7193)

It's pretty straightforward. AED = 90 degrees since AB and CD are perpendicular. So 90 = x + 20, x = 70.